OpenStudy (anonymous):

2) A ball is hit at an angle of 36 degrees to the horizontal from 14 feet above the ground. If it hits the ground 26 feet away (in horizontal distance), what was the initial velocity? Assume there is no air resistance.

OpenStudy (anonymous):

if e=1then let the speed with which ball hits the floor is v so horizontal component of it is vcos36 & vertical vsin36 time taken by ball to come back 2sin36/g so 26feet=(2*vcos36*vsin36)/g so get v as horizontal component is not changed by gravity so initial horizontal component is vcos36 and let vertical component is u so (vsin36)^2-u^2=2*g*14(feets) get u hence the velocity